OpenStudy (anonymous):

select a ordered pair that is a solution of this equation. -2x+6y=-4

OpenStudy (anonymous):

You could select any value for either x or y and evaluate the other one. I'll use x=1 -2(1)+6y=-4 -2+6y=-4 6y=-2 y=-2/6 y=-1/3

OpenStudy (bahrom7893):

are there choices?

OpenStudy (anonymous):

If you don't want fractions, you might have to experiment a bit to find integers.

OpenStudy (bahrom7893):

Like Mathdrills said that's one ordered pair, I could use x=0 and get another one. Actually using x=4 might give u nicer numbers

OpenStudy (anonymous):

yes i have choices (-3,-7) (-7,-3) (-7,6)

OpenStudy (bahrom7893):

okay so try each of those out. For which of these is the equation above true?

OpenStudy (bahrom7893):

That basically means: Is: -2(-2)+6(-7)=-4

OpenStudy (bahrom7893):

Is: -2(-7)+6(-3)=-4

OpenStudy (bahrom7893):

Is: -2(-7)*6(6)=-4?

OpenStudy (bahrom7893):

Do the multiplication/addition and show me which one is true.

OpenStudy (anonymous):

i dont understand ;/ what will the answer be?

OpenStudy (bahrom7893):

Okay this is what it means. An ordered pair must be the solution to that equation. Your choices are: (-3,-7) (-7,-3) (-7,6), or they can also be written as: 1) x=-3; y=-7 2) x=-7; y=-3 3) x=-7; y=6

OpenStudy (bahrom7893):

Plug each of those into the equation: 1) -2x+6y=-4; -2(-3)+6(-7) = -4

OpenStudy (bahrom7893):

Now tell me is this true: -2(-3)+6(-7) = -4 ?

OpenStudy (anonymous):

no ;[

OpenStudy (bahrom7893):

You are right. Moving on to the next point: 2) -2x+6y=-4; -2(-7)+6(-3) = -4; is this one true?

OpenStudy (bahrom7893):

Hello??